how can this be verified csc^2t/cott = csctsect
Answer by Alan3354(69443)

how can this be verified csc^2t/cott = csctsect
-------------
Change to sines and cosines.
(1/sin^2)/(cos/sin) = (1/sin)*(1/cos)
(1/sin)/cos = (1/(sin*cos))
1/sin = 1/sin